# On thermal transpiration and thermomolecular pressure difference

### v1

@@ 2.4. Estimate of $Q_{+}$ and $Q_{-}$
Taking $\beta$ sufficiently small and combining 3.17 with 3.26, we obtain
$\lVert f^{\eta}\rVert^{2}_{L^{2}_{0,\gamma/2}}+$ $\lVert f^{\eta}\rVert^{2}_{L^{2}(\partial\Omega^{+}_{1})}+\lVert(I-P)f^{\eta}\rVert^{2}_{L^{2}(\partial\Omega^{+}_{2})}$ (3.27)
@@
Now, similar to the process used to prove that $f^{\hat{\alpha},\eta}$ is a Cauchy sequence for any fixed $\eta$, we observe that
@@
$\lVert f^{\eta_{1}}-f^{\eta_{2}}\rVert^{2}_{L^{2}_{0,\gamma/2}}+\lVert f^{\eta_{1}}-f^{\eta_{2}}\rVert^{2}_{L^{2}(\partial\Omega^{+}_{1})}+\lVert(I-P)\Big(f^{\eta_{1}}-f^{\eta_{2}}\Big)\rVert^{2}_{L^{2}(\partial\Omega^{+}_{2})}$ (3.29)
$\leq C\max\{\kappa,\kappa^{2}\}\Big(\eta_{1}-\eta_{2}\Big)^{2}\lVert P(f^{\eta_{2}})\rVert^{2}_{L^{2}(\partial\Omega^{-}_{2})}$
@@
This means that $\{f^{\eta}\}_{\eta}$ is a Cauchy sequence in $L^{2}_{0,\gamma/2}$. Therefore $f\coloneqq\lim_{\eta\to 1}f^{\eta}$ is the solution of (3.3). ∎
Chapter 4. $L^{\infty}$ estimate
In this chapter, we denote $w(v)=e^{\alpha|v|^{2}}$. Let $h(x,v)=w(v)f(x,v)$. Then, $h$ satisfies
$\left\{\begin{aligned} v\cdot\nabla_{x}h+\frac{1}{\kappa}L_{w}(h)&=w\phi_{1},&&\quad\text{in}\quad\Omega\times\mathbb{R}^{3},\\ h&=wf_{in},&&\quad\text{on}\quad\partial\Omega^{-}_{1},\\ h&=P_{w}(h)+w\phi_{2}(x,v),&&\quad\text{on}\quad\partial\Omega^{-}_{2},\end{aligned}\right$. (4.1)
@@
by defining $K_{\alpha}(\cdot)\coloneq wK(w^{-1}\cdot)$. Note that $K_{\alpha}$ and $|K_{\alpha}|(f)\coloneq w\int_{\mathbb{R}^{3}}|k(v,u)|w^{-1}f(x,u)du$ have the same properties, where $k(v,u)$ denotes the kernel of $K$. In this chapter, we prove that problem (4.1) admits a unique solution in $L^{\infty}$ when $\Omega$ is unsealed with respect to $\partial\Omega_{1}$, and therefore Theorem 1.1 in section 4.2. First, we consider the approximating problem with the parameters $\hat{a}$ and $\eta$, and we define a recursive sequence $h^{l}$ with $h^{1}=0$.
$\left\{\begin{aligned} \hat{\alpha}\nu(v)h^{l+1}+v\cdot\nabla_{x}h^{l+1}+\frac{1}{\kappa}L_{w}(h^{l+1})&=w\phi_{1},&&\quad\text{in}\quad\Omega\times\mathbb{R}^{3},\\ h^{l+1}&=wf_{in},&&\quad\text{on}\quad\partial\Omega^{-}_{1},\\ h^{l+1}&=\eta P_{w}(h^{l})+w\phi_{2},&&\quad\text{on}\quad\partial\Omega^{-}_{2}.\end{aligned}\right$. (4.2)
Second, we establish the following proposition to relate the $L^{\infty}$ norm of $h^{l}$ to the weighted $L^{2}$ norm of $h^{l}/w$. This relation allows us to apply the results from previous chapter to construct a solution to (4.2) in $L^{\infty}$ and then remove the parameters $\hat{a}$ and $\eta$. This approach is known as the $L^{2}-L^{\infty}$ estimate, which is an useful tool ( [13], [14]) for constructing solutions of Boltzmann equation. Finally we note that proposition 4.1 remains valid when $\hat{a}=0$ and $\eta=1$. The statement of the proposition is
@@
Let $\Omega$ be a bounded set in $\mathbb{R}^{3}$ with $C^{2}$ boundary, and $\Omega$ is completely unsealed with respect to $\partial\Omega_{1}$. Suppose that $h^{l+1}$ is the solution of (4.2) and (1.6) holds. Then, there exists $n$ depending only on the constant of unsealed set $\zeta$ such that
$\lVert h^{l+1}\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}$ (4.3)
@@
$\quad+\frac{1}{8}\max_{0\leq m\leq 2(n-1)}\lVert h^{l-m}\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}+C_{n}\min\{1,\frac{1}{\kappa}\}\max_{0\leq m\leq 2(n-1)}\lVert h^{l+1-m}w^{-1}\rVert_{L^{2}_{0,\gamma/2}(\Omega\times\mathbb{R}^{3})}.$
@@
$|h^{l+1}|(x,v)$ $\leq\frac{1}{\kappa}S^{\kappa}_{\Omega}|K_{\alpha}|(|h^{l+1}|)+S^{\kappa}_{\Omega}(|w\phi_{1}|)+J^{\kappa}\textbf{1}_{\partial\Omega^{-}_{1}}|wf_{in}|+J^{\kappa}\textbf{1}_{\partial\Omega^{-}_{2}}|w\phi_{2}|+J^{\kappa}\textbf{1}_{\partial\Omega^{-}_{2}}P_{w}(|h^{l}|).$ (4.4)
@@
$\displaystyle\int\cdots\int\textbf{1}d\sigma^{1}_{m-1,m}\cdots d\sigma^{1}_{0,1}\leq\zeta\quad\forall x\in\partial\Omega\setminus\partial\Omega_{1},$ (4.6)
@@
Let $\Omega\subset\mathbb{R}^{3}$ be a bounded convex domain with $C^{2}$ boundary and suppose $\Omega$ is unsealed with respect to $\partial\Omega_{1}$. Assume that $3<\gamma\leq 1$, $0\leq a<\frac{1}{4}$ and (1.6) holds. Then, for any $\phi_{1}\in L^{\infty}_{\alpha,-1}(\Omega\times\mathbb{R}^{3}),f_{in}\in L^{\infty}_{\alpha,0}(\partial\Omega_{1}),$ and $\phi_{2}\in L^{\infty}_{\alpha,0}(\partial\Omega_{2})$, there exists a unique solution $h\in L^{\infty}_{\alpha,0}(\Omega\times\mathbb{R}^{3})$ to (4.1). Furthermore. the solution $h$ satisfies following estimate
$\lVert h\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}+\lVert h\rVert_{L^{\infty}(\partial\Omega^{+}_{2})}$ $\leq C_{\zeta}\Big(\lVert f_{in}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{1})}+\lVert\phi_{2}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{2})}+\max\{1,\kappa\}\lVert\phi_{1}\rVert_{L^{\infty}_{a,-1}(\Omega\times\mathbb{R}^{3})}\Big).$ (4.7)
@@
Let $3<\gamma\leq 1$ and assume that (1.6) holds. Then
$|h^{l+1}|(x,v)$ $\leq C_{\zeta}\Big(\lVert f_{in}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{1})}+\lVert\phi_{2}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{2})}+\kappa\lVert\phi_{1}\rVert_{L^{\infty}_{a,-1}(\Omega\times\mathbb{R}^{3})}\Big)$ (4.8)
@@
$\begin{cases}\lVert A_{3}\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}&\leq C\lVert f_{in}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{1})},\\ \lVert A_{4}\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}&\leq C\lVert w(v)\phi_{2}\rVert_{L^{\infty}(\partial\Omega^{-}_{2})}.\end{cases}$ (4.9)
Also, similar to (4.9), we obtain
$\begin{cases}\lVert\sum_{j=1}^{n-1}\hat{P}^{j}(A_{3})\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}&\leq C\frac{1}{1-\zeta}\lVert f_{in}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{1})},\\ \lVert\sum_{j=1}^{n-1}\hat{P}^{j}(A_{4})\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}&\leq C\frac{1}{1-\zeta}\lVert f_{in}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{1})}.\end{cases}$ (4.10)
@@
$\lVert A_{2}\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}$ $\leq C\kappa\lVert\langle v\rangle^{-1}w(v)\phi_{1}\rVert_{L^{\infty}(\partial\Omega^{-}_{1})}$ (4.11)
@@
$\left\lVert\sum_{j=1}^{n-1}\hat{P}^{j}(A_{2})\right\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}$ $\leq C\frac{\kappa}{1-\zeta}\lVert\phi_{1}\rVert_{L^{\infty}_{a,-1}(\Omega\times\mathbb{R}^{3})}.$ (4.12)
@@
$\lVert\hat{P}^{n}(h^{l+1-n})\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}$ $\leq C\zeta^{n-1}\lVert h^{l+1-n}\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}.$ (4.13)
@@
$|h^{l+1}|$ $\leq C\frac{1}{1-\zeta}\Big(\lVert f_{in}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{1})}+\lVert\phi_{2}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{2})}+\kappa\lVert\phi_{1}\rVert_{L^{\infty}_{a,-1}(\Omega\times\mathbb{R}^{3})}\Big)$ (4.14)
@@
$|h^{l+1-j}|$ $\leq C\frac{1}{1-\zeta}\Big(\lVert f_{in}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{1})}+\lVert\phi_{2}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{2})}+\kappa\lVert\phi_{1}\rVert_{L^{\infty}_{\alpha,-1}(\Omega\times\mathbb{R}^{3})}\Big)$ (4.15)
@@
Substitute $h^{l+1-j}$ by 4.15, we obtain
$|h^{l+1}|$ $\leq C_{\zeta}\Big(\lVert f_{in}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{1})}+\lVert\phi_{2}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{2})}+\kappa\lVert\phi_{1}\rVert_{L^{\infty}_{a,-1}(\Omega\times\mathbb{R}^{3})}\Big)$ (4.16)
@@
$\quad+\frac{\zeta^{n-1}}{1-\zeta}\max_{0\leq j\leq n-1}\lVert h^{l+1-j-n}\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}+\left(\frac{1}{\kappa}\right)^{2}\sum_{j=0}^{n-1}\sum_{i=0}^{n-1}\hat{P}^{j}S^{\kappa}_{\Omega}|K_{\alpha}|\hat{P}^{i}S^{\kappa}_{\Omega}|K_{\alpha}|(|h^{l+1-j-i}|).$ (4.17)
@@
$A_{ij}(f)\coloneq\left(\frac{1}{\kappa}\right)^{2}\hat{P}^{j}S^{\kappa}_{\Omega}|K_{\alpha}|\hat{P}^{i}S^{\kappa}_{\Omega}|K_{\alpha}|(f)$ (4.18)
@@
Suppose $\Omega$ is a bounded set in $\mathbb{R}^{3}$ with $C^{2}$ boundary and $\Omega$ is completely unsealed with respect to $\partial\Omega_{1}$. Given $\epsilon>0$ and $i,j\in\mathbb{N}$. Suppose $3<\gamma\leq 1$ and (1.6) holds. Then
$\lVert A_{ij}(f)\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}\leq C\zeta^{j+i-2}\Big(\epsilon\lVert f\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}+C_{\epsilon}\min\{1,\frac{1}{\kappa}\}\lVert f/w\rVert_{L^{2}_{0,\gamma/2}(\Omega\times\mathbb{R}^{3})}\Big)$ (4.19)
@@
Without loss of generality, we assume $f\geq 0$. For $i=j=0$, we use properties of $|k_{\alpha}|$ like 2.3, proposition 2.6 , and proposition 2.5. By definition,
$A_{00}$ $=\left(\frac{1}{\kappa}\right)^{2}\int_{0}^{\tau_{-}(x,v)}\int_{\mathbb{R}^{3}}\int_{0}^{\tau_{-}(x-tv,u)}\int_{\mathbb{R}^{3}}e^{-\frac{1}{\kappa}\nu(v)t}|k_{\alpha}|(v,u)e^{-\frac{1}{\kappa}\nu(u)s}|k_{\alpha}|(u,w)f(x-tv-su,w)dwdsdudt.$ (4.20)
@@
$A=\{|v-u|>\epsilon_{2},|u|<N\}\text{ and }B=\{|u-w|>\epsilon_{1}\}.$ (4.21)
@@
$+\left(\frac{1}{\kappa}\right)^{2}\int_{0}^{\tau_{-}(x,v)}\int_{A}\int_{1/N}^{\tau_{-}(x-tv,u)}\int_{B}e^{-\frac{1}{\kappa}\nu(v)t}|k_{\alpha}|(v,u)e^{-\frac{1}{\kappa}\nu(u)s}|k_{\alpha}|(u,w)f(x-tv-su,w)dwdsdudt.$ (4.23)
@@
$\quad+C\left(\frac{1}{\kappa}\right)^{2}\textbf{1}_{|v|<M}\Big(\int_{\frac{\mathcal{D}}{5M}}^{\frac{\mathcal{D}}{|v|}}\int_{\mathbb{R}^{3}}\int_{1/N}^{\infty}\int_{\mathbb{R}^{3}}e^{-\frac{1}{\kappa}\nu(v)t}\frac{1}{s^{3}}\nu(w)f^{2}(y,w)w^{-2}(w)dwdsdydt\Big)^{1/2}$
$\leq C\Big(\epsilon^{3-|\gamma|}_{1}+\epsilon^{3-|\gamma|}_{2}+\frac{1}{1+N}+\frac{1}{N}+\frac{1}{M}\Big)\lVert f\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}+C\left(\frac{1}{\kappa}\right)\lVert f/w\rVert_{L^{2}_{0,\gamma/2}(\Omega\times\mathbb{R}^{3})}.$ (4.24)
@@
$\hat{P}S^{\kappa}_{\Omega}|K_{\alpha}|(f)(x,v)$ (4.25)
@@
$\hat{P}S^{\kappa}_{\Omega}|K_{\alpha}|(f)(x,v)$ (4.27)
@@
$\quad+\frac{1}{\kappa}\displaystyle\int_{\{v_{1}\cdot n(q)>0\}\cap A}\int_{0}^{\frac{\mathcal{D}}{|v_{1}|}}\int_{B}e^{-\frac{1}{\kappa}\nu(v_{1})t}|k|(v_{1},u)f(q-tv_{1},u)w^{-1}(u)m^{1/2}(v_{1})v_{1}\cdot n(q)dudtdv_{1}.$ (4.28)

### v2

@@ 2.4. Estimate of $Q_{+}$ and $Q_{-}$
Taking $\beta$ sufficiently small and combining (66) with (75), we obtain
$\lVert f^{\eta}\rVert^{2}_{L^{2}_{0,\gamma/2}}+$ $\lVert f^{\eta}\rVert^{2}_{L^{2}(\partial\Omega^{+}_{1})}+\lVert(I-P)f^{\eta}\rVert^{2}_{L^{2}(\partial\Omega^{+}_{2})}$ (76)
@@
We observe that
@@
$\lVert f^{\eta_{1}}-f^{\eta_{2}}\rVert^{2}_{L^{2}_{0,\gamma/2}}+\lVert f^{\eta_{1}}-f^{\eta_{2}}\rVert^{2}_{L^{2}(\partial\Omega^{+}_{1})}+\lVert(I-P)\Big(f^{\eta_{1}}-f^{\eta_{2}}\Big)\rVert^{2}_{L^{2}(\partial\Omega^{+}_{2})}$ (78)
$\leq C\max\{\kappa^{-2},\kappa^{2}\}\Big(\eta_{1}-\eta_{2}\Big)^{2}\lVert P(f^{\eta_{2}})\rVert^{2}_{L^{2}(\partial\Omega^{-}_{2})}$
@@
This means that $\{f^{\eta}\}_{\eta}$ is a Cauchy sequence in $L^{2}_{0,\gamma/2}$. Therefore $f\coloneqq\lim_{\eta\to 1}f^{\eta}$ is the solution of (52). ∎
Section 4. $L^{\infty}$ estimate
In this section, we denote $w(v)=e^{\alpha|v|^{2}}$. Let $h(x,v)=w(v)f(x,v)$. Then, $h$ satisfies
$\left\{\begin{aligned} v\cdot\nabla_{x}h+\frac{1}{\kappa}L_{w}(h)&=w\phi_{1},&&\quad\text{in}\quad\Omega\times\mathbb{R}^{3},\\ h&=wf_{in},&&\quad\text{on}\quad\partial\Omega^{-}_{1},\\ h&=P_{w}(h)+w\phi_{2}(x,v),&&\quad\text{on}\quad\partial\Omega^{-}_{2},\end{aligned}\right$. (80)
@@
by defining $K_{\alpha}(\cdot)\coloneq wK(w^{-1}\cdot)$. Note that $K_{\alpha}$ and $|K_{\alpha}|(f)\coloneq w\int_{\mathbb{R}^{3}}|k(v,u)|w^{-1}f(x,u)du$ have the same properties, where $k(v,u)$ denotes the kernel of $K$. In this section, we prove that problem (80) admits a unique solution in $L^{\infty}$ when $\Omega$ is unsealed with respect to $\partial\Omega_{1}$, and therefore Theorem 1.1 in subsection 4.2. First, we consider the approximating problem with the parameters $\hat{a}$ and $\eta$, and we define a recursive sequence $h^{l}$ with $h^{1}=0$.
$\left\{\begin{aligned} \hat{\alpha}\nu(v)h^{l+1}+v\cdot\nabla_{x}h^{l+1}+\frac{1}{\kappa}L_{w}(h^{l+1})&=w\phi_{1},&&\quad\text{in}\quad\Omega\times\mathbb{R}^{3},\\ h^{l+1}&=wf_{in},&&\quad\text{on}\quad\partial\Omega^{-}_{1},\\ h^{l+1}&=\eta P_{w}(h^{l})+w\phi_{2},&&\quad\text{on}\quad\partial\Omega^{-}_{2}.\end{aligned}\right$. (81)
Second, we establish the following proposition to relate the $L^{\infty}$ norm of $h^{l}$ to the weighted $L^{2}$ norm of $h^{l}/w$. This relation allows us to apply the results from previous section to construct a solution to (81) in $L^{\infty}$ and then remove the parameters $\hat{a}$ and $\eta$. This approach is known as the $L^{2}-L^{\infty}$ estimate, which is an useful tool ( [13], [14]) for constructing solutions of Boltzmann equation. Finally we note that proposition 4.1 remains valid when $\hat{a}=0$ and $\eta=1$. The statement of the proposition is
@@
Let $\Omega$ be a bounded set in $\mathbb{R}^{3}$ with $C^{2}$ boundary, and $\Omega$ is completely unsealed with respect to $\partial\Omega_{1}$. Suppose that $h^{l+1}$ is the solution of (81) and (8) holds. Then, there exists $n$ depending only on the constant of unsealed set $\zeta$ such that
$\lVert h^{l+1}\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}$ (82)
@@
$\quad+\frac{1}{8}\max_{0\leq m\leq 2(n-1)}\lVert h^{l-m}\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}+\frac{C_{\zeta}}{\kappa}\max_{0\leq m\leq 2(n-1)}\lVert h^{l+1-m}w^{-1}\rVert_{L^{2}_{0,\gamma/2}(\Omega\times\mathbb{R}^{3})}.$
for any $l\geq 2n$.
@@
$|h^{l+1}|(x,v)$ $\leq\frac{1}{\kappa}S^{\kappa}_{\Omega}|K_{\alpha}|(|h^{l+1}|)+S^{\kappa}_{\Omega}(|w\phi_{1}|)+J^{\kappa}\textbf{1}_{\partial\Omega^{-}_{1}}|wf_{in}|+J^{\kappa}\textbf{1}_{\partial\Omega^{-}_{2}}|w\phi_{2}|+J^{\kappa}\textbf{1}_{\partial\Omega^{-}_{2}}P_{w}(|h^{l}|).$ (83)
@@ 4.1. The existence of solution to (80)
$\displaystyle\int\cdots\int\textbf{1}d\sigma^{1}_{m-1,m}\cdots d\sigma^{1}_{0,1}\leq\zeta\quad\forall x\in\partial\Omega\setminus\partial\Omega_{1},$ (85)
@@
Let $\Omega\subset\mathbb{R}^{3}$ be a bounded convex domain with $C^{2}$ boundary and suppose $\Omega$ is unsealed with respect to $\partial\Omega_{1}$. Assume that $3<\gamma\leq 1$, $0\leq a<\frac{1}{4}$ and (8) holds. Then, for any $\phi_{1}\in L^{\infty}_{\alpha,-1}(\Omega\times\mathbb{R}^{3}),f_{in}\in L^{\infty}_{\alpha,0}(\partial\Omega_{1}),$ and $\phi_{2}\in L^{\infty}_{\alpha,0}(\partial\Omega_{2})$, there exists a unique solution $h\in L^{\infty}_{\alpha,0}(\Omega\times\mathbb{R}^{3})$ to (80). Furthermore. the solution $h$ satisfies following estimate
$\lVert h\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}+\lVert h\rVert_{L^{\infty}(\partial\Omega^{+}_{2})}$ (86)
$\leq C_{\zeta}\max\{1,\kappa^{-1}\}\Big(\lVert f_{in}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{1})}+\lVert\phi_{2}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{2})}+\max\{1,\kappa\}\lVert\phi_{1}\rVert_{L^{\infty}_{a,-1}(\Omega\times\mathbb{R}^{3})}\Big).$
@@
Let $3<\gamma\leq 1$ and assume that (8) holds. Then
$|h^{l+1}|(x,v)$ $\leq C_{\zeta}\Big(\lVert f_{in}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{1})}+\lVert\phi_{2}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{2})}+\kappa\lVert\phi_{1}\rVert_{L^{\infty}_{a,-1}(\Omega\times\mathbb{R}^{3})}\Big)$ (87)
@@
$\begin{cases}\lVert A_{3}\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}&\leq C\lVert f_{in}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{1})},\\ \lVert A_{4}\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}&\leq C\lVert w(v)\phi_{2}\rVert_{L^{\infty}(\partial\Omega^{-}_{2})}.\end{cases}$ (88)
Also, similar to (88), we obtain
$\begin{cases}\lVert\sum_{j=1}^{n-1}\hat{P}^{j}(A_{3})\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}&\leq C\frac{1}{1-\zeta}\lVert f_{in}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{1})},\\ \lVert\sum_{j=1}^{n-1}\hat{P}^{j}(A_{4})\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}&\leq C\frac{1}{1-\zeta}\lVert f_{in}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{1})}.\end{cases}$ (89)
@@
$\lVert A_{2}\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}$ $\leq C\kappa\lVert\langle v\rangle^{-1}w(v)\phi_{1}\rVert_{L^{\infty}(\partial\Omega^{-}_{1})}$ (90)
@@
$\left\lVert\sum_{j=1}^{n-1}\hat{P}^{j}(A_{2})\right\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}$ $\leq C\frac{\kappa}{1-\zeta}\lVert\phi_{1}\rVert_{L^{\infty}_{a,-1}(\Omega\times\mathbb{R}^{3})}.$ (91)
@@
$\lVert\hat{P}^{n}(h^{l+1-n})\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}$ $\leq C\zeta^{n-1}\lVert h^{l+1-n}\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}.$ (92)
@@
$|h^{l+1}|$ $\leq C\frac{1}{1-\zeta}\Big(\lVert f_{in}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{1})}+\lVert\phi_{2}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{2})}+\kappa\lVert\phi_{1}\rVert_{L^{\infty}_{a,-1}(\Omega\times\mathbb{R}^{3})}\Big)$ (93)
@@
$|h^{l+1-j}|$ $\leq C\frac{1}{1-\zeta}\Big(\lVert f_{in}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{1})}+\lVert\phi_{2}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{2})}+\kappa\lVert\phi_{1}\rVert_{L^{\infty}_{\alpha,-1}(\Omega\times\mathbb{R}^{3})}\Big)$ (94)
@@
Substitute $h^{l+1-j}$ by 94, we obtain
$|h^{l+1}|$ $\leq C_{\zeta}\Big(\lVert f_{in}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{1})}+\lVert\phi_{2}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{2})}+\kappa\lVert\phi_{1}\rVert_{L^{\infty}_{a,-1}(\Omega\times\mathbb{R}^{3})}\Big)$ (95)
@@
$\quad+\frac{\zeta^{n-1}}{1-\zeta}\max_{0\leq j\leq n-1}\lVert h^{l+1-j-n}\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}+\left(\frac{1}{\kappa}\right)^{2}\sum_{j=0}^{n-1}\sum_{i=0}^{n-1}\hat{P}^{j}S^{\kappa}_{\Omega}|K_{\alpha}|\hat{P}^{i}S^{\kappa}_{\Omega}|K_{\alpha}|(|h^{l+1-j-i}|).$ (96)
@@
$A_{ij}(f)\coloneq\left(\frac{1}{\kappa}\right)^{2}\hat{P}^{j}S^{\kappa}_{\Omega}|K_{\alpha}|\hat{P}^{i}S^{\kappa}_{\Omega}|K_{\alpha}|(f)$ (97)
@@
Suppose $\Omega$ is a bounded set in $\mathbb{R}^{3}$ with $C^{2}$ boundary and $\Omega$ is completely unsealed with respect to $\partial\Omega_{1}$. Given $\epsilon>0$ and $i,j\in\mathbb{N}$. Suppose $3<\gamma\leq 1$ and (8) holds. Then
$\lVert A_{ij}(f)\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}\leq C\zeta^{j+i-2}\Big(\epsilon\lVert f\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}+C_{\epsilon}\min\{1,\frac{1}{\kappa}\}\lVert f/w\rVert_{L^{2}_{0,\gamma/2}(\Omega\times\mathbb{R}^{3})}\Big)$ (98)
@@
Without loss of generality, we assume $f\geq 0$. For $i=j=0$, we use properties of $|k_{\alpha}|$ like 29, proposition 2.6 , and proposition 2.5. By definition,
$A_{00}$ $=\left(\frac{1}{\kappa}\right)^{2}\int_{0}^{\tau_{-}(x,v)}\int_{\mathbb{R}^{3}}\int_{0}^{\tau_{-}(x-tv,u)}\int_{\mathbb{R}^{3}}e^{-\frac{1}{\kappa}\nu(v)t}|k_{\alpha}|(v,u)e^{-\frac{1}{\kappa}\nu(u)s}|k_{\alpha}|(u,w)f(x-tv-su,w)dwdsdudt.$ (99)
@@
$A=\{|v-u|>\epsilon_{2},|u|<N\}\text{ and }B=\{|u-w|>\epsilon_{1}\}.$ (100)
@@
$+\left(\frac{1}{\kappa}\right)^{2}\int_{0}^{\tau_{-}(x,v)}\int_{A}\int_{1/N}^{\tau_{-}(x-tv,u)}\int_{B}e^{-\frac{1}{\kappa}\nu(v)t}|k_{\alpha}|(v,u)e^{-\frac{1}{\kappa}\nu(u)s}|k_{\alpha}|(u,w)f(x-tv-su,w)dwdsdudt.$ (102)
@@
$\quad+C\left(\frac{1}{\kappa}\right)^{2}\textbf{1}_{|v|<M}\Big(\int_{\frac{\mathcal{D}}{5M}}^{\frac{\mathcal{D}}{|v|}}\int_{\mathbb{R}^{3}}\int_{1/N}^{\infty}\int_{\mathbb{R}^{3}}e^{-\frac{C_{M}}{\kappa}t}\frac{1}{s^{3}}\nu(w)f^{2}(y,w)w^{-2}(w)dwdsdydt\Big)^{1/2}$
$\leq C\Big(\epsilon^{3-|\gamma|}_{1}+\epsilon^{3-|\gamma|}_{2}+\frac{1}{1+N}+\frac{1}{N}+\frac{1}{M}\Big)\lVert f\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}+C\min\{1,\frac{1}{\kappa}\}\lVert f/w\rVert_{L^{2}_{0,\gamma/2}(\Omega\times\mathbb{R}^{3})}.$ (103)
@@
$\hat{P}S^{\kappa}_{\Omega}|K_{\alpha}|(f)(x,v)$ (104)
@@
$\hat{P}S^{\kappa}_{\Omega}|K_{\alpha}|(f)(x,v)$ (106)
@@
$\quad+\frac{1}{\kappa}\displaystyle\int_{\{v_{1}\cdot n(q)>0\}\cap A}\int_{0}^{\frac{\mathcal{D}}{|v_{1}|}}\int_{B}e^{-\frac{1}{\kappa}\nu(v_{1})t}|k|(v_{1},u)f(q-tv_{1},u)w^{-1}(u)m^{1/2}(v_{1})v_{1}\cdot n(q)dudtdv_{1}.$ (107)